OpenStudy (anonymous):

Again just checking my answer but don't know if it would be 16 or -16?

OpenStudy (anonymous):

.

OpenStudy (anonymous):

\(g(g(x)) = -(-x^2)^2 = -(x^2)^2 = -x^4\)

OpenStudy (anonymous):

So, \(g(g(2))=...?\)

OpenStudy (usukidoll):

g o g (x) means put the g(x) function within itself \[g(x) = -x^2 \] \[(g o g) (x)= -(-x^2)^2\] expanding that out gives \[(g o g) (x)= -(-x^2)(-x^2)\] while a negative times a negative is a positive, there is still a negative sign hanging outside \[(g o g) (x)= -(x^4)\] when we evaluate the function at x = 2 \[(g o g) (x)= -(2^4) = -16\]